OpenStudy (anonymous):

What is the sum of the angle measures of a 37-gon?

OpenStudy (anonymous):

There is an equation to solve this. It looks like: Sum of interior angles = (n - 2) * 180 degrees where n is the number of sides of the n-gon. In this case, 37.

OpenStudy (anonymous):

the sum of all angles is (2n-4)*90 here n=37

OpenStudy (anonymous):

(2n-4)*90 =(2*37-4)*90=(74-4)*90=70*90=6300

OpenStudy (anonymous):

Sum of Interior angles: (3n-6)*60

OpenStudy (anonymous):

In all seriousness you should use Logic's simplified formula.

OpenStudy (tkhunny):

Is it a REGULAR 37-gon? Does it matter? I like to remember that the sum of EXTERNAL angles is ALWAYS 360º. \(360º / 37 = 9.\overline{729}º\) -- The measure of each External Angle. Internal Angels and External Angles are Supplementary \(180º - 9.\overline{729}º = 170.\overline{270}º\) -- The measure of each Internal Angle. I don't actually recommend this method over the \((n-2)\cdot 180º\) formula for such a question as the sum of the internal angles, but it can be useful if you forget.

OpenStudy (anonymous):

The sum of the interior angles of a convex n-gon is (n - 2) * 180 Notice that I assume we are dealing with a convex polygon, as a concave one will wield infinitely many possible solutions. Since we have a 37-gon here, the sum of the interior angles will be: (37 - 2) * 180 = 6300 degrees